OpenStudy (anonymous):

Write 3 ^2 x 3 ^-3 x 3 ^6 as a single exponent. 3 ^4 3 ^5 3 ^11 27 ^5

OpenStudy (cj49):

3^2x3^6/3^3=?

OpenStudy (anonymous):

243

OpenStudy (cj49):

correct 3^2=3x3 3^6=3x3x3x3x3x3 3x3x3x3x3x3x3x3/3x3x3 =3^5=243

OpenStudy (anonymous):

C: